OpenStudy (anonymous):

Write down the matrix a=[aij] where aij=jsin(ipi/4)

OpenStudy (anonymous):

\[A=[aij] -> aij=j \sin (\frac{ i \pi }{ 4 })\]

OpenStudy (anonymous):

I have no idea how to do this so if you could show me

OpenStudy (anonymous):

\[A=\left[\begin{matrix}a_{i,j}\end{matrix}\right], \text{ where }a_{i,j}=j\sin\left(\frac{i\pi}{4}\right)\] Depending on the dimension of A, you'll have a matrix where any given element in row (i) and column (j) is of the form given (j sin(i pi/4)). This means, for example, that the second row, third column is occupied by \[a_{2,3}=3\sin\left(\frac{2\pi}{4}\right)=3\] Suppose A is a 2x2 matrix. Then you have \[\left[\begin{matrix}a_{1,1}&a_{1,2}\\a_{2,1}&a_{2,2}\end{matrix}\right]=\left[\begin{matrix}\sin\left(\frac{\pi}{4}\right)&2\sin\left(\frac{\pi}{4}\right)\\\sin\left(\frac{2\pi}{4}\right)&2\sin\left(\frac{2\pi}{4}\right)\end{matrix}\right]=\left[\begin{matrix}\frac{1}{\sqrt2}&\frac{2}{\sqrt2}\\1&2\end{matrix}\right]\]

OpenStudy (anonymous):

Oh I must have missed notes on this stuff.Do you just assume that it is a 2x2 matrix?

OpenStudy (anonymous):

Not necessarily, I just used that as an example.

OpenStudy (anonymous):

Oh okay thanks!!